# Breaking Down and Building Up: Mixture of Skill-Based Vision-and-Language Navigation Agents

### v1

@@ Abstract
Vision-and-Language Navigation (VLN) poses significant challenges in enabling agents to interpret natural language instructions and navigate complex 3D environments. While recent progress has been driven by large-scale pre-training and data augmentation, current methods still struggle to generalize to unseen scenarios, particularly when complex spatial and temporal reasoning is required. In this work, we propose SkillNav, a modular framework that introduces structured, skill-based reasoning into Transformer-based VLN agents. Our method decomposes navigation into a set of interpretable atomic skills (e.g., Vertical Movement, Area and Region Identification, Stop and Pause), each handled by a specialized agent. We then introduce a novel zero-shot Vision-Language Model (VLM)-based router, which dynamically selects the most suitable agent at each time step by aligning sub-goals with visual observations and historical actions. SkillNav achieves a new state-of-the-art performance on the R2R benchmark and demonstrates strong generalization to the GSA-R2R benchmark that includes novel instruction styles and unseen environments.
Codehttps://github.com/HLR/SkillNav
@@ Introduction
Vision-and-Language Navigation (VLN) (Anderson et al. 2018b; Zhang et al. 2024c) is a critical subfield of embodied AI that integrates natural language understanding, visual perception, and sequential decision-making to allow autonomous agents to navigate and interact within visual environments. With the rise of foundation models e.g., Large Language Models (LLM) and Vision-Language Models (VLMs) (Zhou et al. 2024a; Xiao and Zhu 2025; Li et al. 2024; Zhang et al. 2024a, b), VLN has seen notable progress in multi-modal grounding and generalization.
@@
Figure 1: SkillNav decomposes complex navigation instructions into atomic skills, which can then be flexibly recomposed to address new environments featuring diverse instruction styles and visual scenarios.
@@
Despite recent advances, a key challenge in VLN lies in enabling agents to generalize reliably and interact with unseen environments and novel instructions. Previous approaches have enhanced VLN agents’ generalization ability through extensive data augmentation (Hao et al. 2020a; Chen et al. 2022a; Wang et al. 2023, 2024b) and training agents on large-scale synthetic instruction-trajectory pairs across varied environments. While data-driven methods improve VLN agents’ generalization, their main limitation is reliance on black-box, end-to-end models (Anderson et al. 2018b; Hong et al. 2021) that tend to memorize training examples. This restricts their effectiveness in unobserved scenarios requiring deeper compositional reasoning, such as understanding diverse instructions, temporal relationships, or complex landmarks, and generalizing across a wide range of visual environments.
@@
To address these limitations, we propose SkillNav, a modular VLN framework that decomposes navigation learning into individual and reusable skills, enabling flexible re-composition and enhanced generalization in new environments (as shown in Fig 1). Unlike prior methods that treat instruction execution as an end-to-end mapping from instructions directly to actions, SkillNav explicitly captures the compositional nature of navigation tasks. Furthermore, we introduce a novel VLM-based router that leverages multi-modal reasoning to dynamically select the most appropriate skill at each navigation step, conditioned on the current sub-instruction, visual observation, and historical actions. SkillNav not only improves interpretability by making the decision-making processes more transparent but also facilitates robust adaptation to diverse instructions and unseen visual environments.
We identify a set of atomic skills required for effectively completing the VLN task. For each skill, we curate a dataset containing relevant instructions paired with corresponding visual observations. Each skill-based agent, built upon a strong VLN backbone, is then independently fine-tuned using its specific skill-specific dataset.
@@
Moreover, we introduce a temporal reordering module to generate chronologically ordered sub-instructions, facilitating effective temporal reasoning during skill selection. Finally, we integrate a VLM-based router that dynamically identifies the next relevant sub-instruction and selects the most suitable skill-based agent to execute the corresponding navigation action.
@@
SkillNav achieves state-of-the-art (SOTA) performance on the Room-to-Room (R2R) benchmark (Anderson et al. 2018b), and the GSA-R2R dataset (Hong et al. 2025) which introduces novel instructions and diverse visual environments, including both unseen residential and non-residential settings. Additionally, we evaluate individual skill-based agents using NavNuances (Wang et al. 2024c), a dataset specifically designed for fine-grained skill evaluation. We provide comprehensive ablation study and qualitative analysis to thoroughly assess the effectiveness of each component within our framework and justify our router design choices.
Our contributions are summarized as follows:
@@
    We propose SkillNav, a modular framework that explicitly decomposes navigation tasks into atomic, reusable skills, significantly enhancing generalization to novel instructions and visual environments.
@@
    We introduce a novel VLM-based router that dynamically selects the most suitable skill at each navigation step by aligning sub-goals with visual observations and historical actions.
@@
    We achieve SOTA performance on the R2R benchmark and demonstrate robust generalization on the challenging GSA-R2R dataset and provide comprehensive analysis and ablation study.
@@
Vision-and-Language Navigation Models.
A wide range of methods have been proposed for addressing VLN tasks. These methods have evolved from early LSTM-based architectures (Anderson et al. 2018b; Tan, Yu, and Bansal 2019) to Transformer-based models (Chen et al. 2021, 2022b; An et al. 2023) and, most recently, to Large Language Model (LLM)-based agents (Zhou, Hong, and Wu 2023; Chen et al. 2024; Lin et al. 2024; Zhou et al. 2024b; Zheng et al. 2024; Zhang et al. 2025). A central and ongoing challenge in VLN research is enhancing the generalization capability of agents, allowing them to navigate effectively in unfamiliar environments and handle novel instructions. To enhance generalization, most existing methods utilize data-driven augmentation strategies, focusing either on augmenting visual observations (Li, Tan, and Bansal 2022; Liu et al. 2021; Li and Bansal 2023) or synthesizing additional navigation instructions (Wang et al. 2023, 2024b; Hao et al. 2020b; Zhang and Kordjamshidi 2023; Zhang, Guo, and Kordjamshidi 2024). However, a fundamental limitation of purely data-driven augmentation approaches lies in their reliance on end-to-end training paradigms. Such monolithic models often memorize training examples rather than genuinely generalize, failing to fundamentally address the compositional reasoning required by novel or unseen scenarios. In contrast, our work directly addresses this limitation by introducing a modular framework that decomposes the VLN task into a set of reusable navigation skills. This modular design allows flexible recomposition of learned skills, substantially improving generalization to novel environments and diverse instruction types. Additionally, our approach enhances interpretability by transparently revealing internal decision-making processes.
Skill-based MoE Systems.
Mixture-of-Experts (MoE) models traditionally operate at the parameter level, distributing input across multiple expert networks to improve capacity and efficiency (Jacobs et al. 1991; Jordan and Jacobs 1994; Yuksel, Wilson, and Gader 2012). Sparsely activated MoEs (Shazeer et al. 2017; Lepikhin et al. 2021; Zhang et al. 2021; Zuo et al. 2022) further scale this idea by routing each input to a small subset of experts, making it possible to train trillion-parameter models while controlling inference cost. More recently, large language models have begun to employ skill-based MoEs at the module or LLM level, where different LLMs are specialized through fine-tuning or task profiling (Riquelme et al. 2021; Wang et al. 2024a; Dai et al. 2024; Jiang et al. 2024; Xue et al. 2024; Chen et al. 2025a; Yu et al. 2025; Zhou et al. 2024c), and expert selection is performed via prompting or routing mechanisms based on task semantics. While these methods significantly enhance modularity and interpretability, they have not yet been explored within the VLN domain. Our work introduces a novel skill-based MoE framework specifically designed for VLN tasks, which employs a set of specialized agents to improve generalization and interpretability.
Preliminaries
In the VLN task, an agent navigates through an environment by following a natural language instruction $I$ to reach a specified target location. The environment is discretized into a connectivity graph $\mathcal{G}=(V,E)$, where $V$ denotes a non-empty set of navigable nodes, and $E$ is a set of undirected connectivity edges. At each time step $t$, the agent located at viewpoint $v_{t}$ receives a panorama represented by $n$ images, denoted as $D_{t}=\{o_{i}\}_{i=0}^{n}$. The agent is aware of a subset of views $O_{t}\subseteq D_{t}$ heading towards its navigable neighboring nodes $\mathcal{N}(v_{t})$. The local action space $A_{t}$ contains navigating to node $v\in\mathcal{N}(v_{t})$ or stopping at current node $v_{t}$.
$a^{*}_{t}=\pi(I,O_{t},M_{t}),$ (1)
where $M_{t}\subseteq\mathcal{G}$ represents online constructed topological map that has been observed after $t$ steps of navigation, and $a^{*}_{t}\in A_{t}$ represents the predicted action.
Methodology
We propose a framework for VLN that coordinates a set of atomic skill-based agents to solve navigation tasks. As shown in Figure 2, the framework comprises three components: a temporal reordering module for instruction decomposition, a VLM-based router for skill selection, and a set of skill-specific agents. Each agent is built upon the DUET architecture and trained with tailored synthetic data to make skill-conditioned decisions. This section introduces the proposed skill taxonomy, skill-specific synthetic dataset construction, and reasoning framework for acquiring these modular skills.
@@
The atomic skills, as defined in NavNuances (Wang et al. 2024c), are considered essential for building a robust VLN agent. In this work, we adopt four frequently observed atomic skills from NavNuances, Direction Adjustment, Vertical Movement, Landmark Detection, and Area and Region Identification.
However, we find persistent challenges in temporal reasoning and stop criteria. Errors in temporal reasoning often disrupt the correct order of subgoal execution. Stop decisions are sometimes made too early or too late, reducing navigation success. To address these issues, we extend the skill taxonomy with two additional skills: Stop and Pause and Temporal Order Planning.
In the following, we elaborate on these two new skills and their roles in navigation.
  1. 1.
    Stop and Pause captures the agent’s ability to dynamically control motion termination and temporary halting in response to visual or linguistic cues. This includes recognizing explicit stop commands (e.g., “Stop at the doorway”) or context-sensitive halts triggered by landmarks or obstacles (e.g., “Pause when you see the red sign”). Unlike generic action execution, stop and pause skills emphasize precise temporal-spatial control to ensure safe, context-aware navigation.
  2. 2.
    Temporal Order Planning captures the agent’s ability to reason over the sequence and structure of subgoals. This includes understanding conditional immediacy (e.g., “Once you enter the hallway, turn left”), maintaining actions for a bounded duration (e.g., “Keep walking until you see the staircase”), executing forward sequential steps (e.g., “Go forward, then turn right, and finally stop”), and handling backward references to prior states (e.g., “Before turning, make sure you’re at the hallway entrance”). Unlike low-level action chaining, temporal order planning captures higher-level temporal logic that guides when and how atomic skills should be executed in order.
To quantify the presence and frequency of these skills in R2R (Anderson et al. 2018a), we perform a keyword-based analysis of the navigation instructions as shown in Figure 3 following the skill inference in Symbolic-MoE (Chen et al. 2025b). Each instruction is scanned for a curated set of indicative keywords, compiled for each skill category based on linguistic patterns observed in prior datasets and real-world navigation discourse. For instance, terms like “after” or “next” are used to detect Temporal Order Planning, while words such as “stairs” or “elevator” signal Vertical Movement. An instruction can be counted for multiple skills if it exhibits multiple relevant keywords.
@@
To enable the training of skill-specialized agents, we construct a set of synthetic datasets in which each trajectory–instruction pair is intentionally designed to emphasize a single navigation skill.
We begin with a random starting node in the Matterport3D (Chang et al. 2017) environment and sample diverse navigation paths through graph traversal. For each skill, we define filtering heuristics to select trajectories where this skill is the primary factor for successful navigation. For instance, we keep paths containing significant elevation changes for the Vertical Movement skill. And for the Direction Adjustment category, we emphasize frequent orientation changes or non-trivial turning sequences. Each selected trajectory consists of a sequence of panoramic observations. We constrain trajectory length to 4–7 steps to keep the difficulty and temporal context comparable to natural VLN data. Additional details are provided in the supplementary material.
To generate skill-focused instruction, we feed the observation sequence of each candidate trajectory into GPT-4o (OpenAI 2024) with a structured prompt. We design the prompts such that the generated instructions preserve the general linguistic quality of real VLN datasets, including comparable sentence length, vocabulary diversity, and fluency, while intentionally emphasizing the content toward the targeted skill. This is achieved by providing GPT-4o with explicit skill-focused cues during generation, encouraging, for example, frequent references to orientation change for the Direction Adjustment skill or strong emphasis on landmark description for the Landmark Detection skill. For each skill, we synthesize $N$ such trajectory–instruction pairs, forming six separate datasets. A summary of dataset statistics is provided in Table 1.
The training of each skill-based agent is conducted in two stages. In the first stage, we fine-tune the pre-trained DUET model using the R2R training split, the ScaleVLN augmentation data (Wang et al. 2023), and our Temporal Synthetic dataset to obtain a strong, skill-agnostic backbone. In the second stage, this backbone is further fine-tuned on another skill-specific synthetic dataset, enabling the agent to specialize in the targeted skill. Following this process, we obtain five specialized skill-based agents: the Direction Adjustment agent ($\pi_{\text{da}}$), Vertical Movement agent ($\pi_{\text{vm}}$), Stop and Pause agent ($\pi_{\text{sp}}$), Landmark Detection agent ($\pi_{\text{ld}}$), and Area and Region Identification agent ($\pi_{\text{ar}}$).
@@
After training specialized agents for different navigation sub-tasks, we introduce our SkillNav framework in this section. Specifically, SkillNav first employs a temporal reordering module to generate structured execution plans. Then, we introduce a VLM-based action router to accurately identify the current subgoal and dynamically select the corresponding skill-based agent to execute the appropriate action.
@@
Our Mixture of Skill-based VLN
SkillNav (ours) 11 $1.97$ $\mathbf{89}$ $83$ $77$ $2.53$ $83$ $78$ $70$ $\mathbf{79}$ $\mathbf{69}$ $\mathbf{72}$ $\mathbf{61}$ $\mathbf{57}$ $\mathbf{48}$
Table 2: Performance comparison on R2R and GSA-R2R benchmarks. indicates large-scale data augmentation. SRDF performs best on R2R due to extensive pretraining on data that mimics R2R-style instructions; however, it struggles to generalize effectively to the more challenging GSA-R2R dataset.
@@
The Temporal Order Module only takes the original natural language instruction as input. It applies the instruction reordering prompt to turn navigation instructions into subgoals $I_{\text{reorder}}$. It follows the four temporal logics described in the Skill Taxonomy section, making implicit temporal details explicit and ensuring the correct subgoal execution order. This procedure is formulated as
@@
To coordinate skill-based agents during navigation, we introduce an Action Router that dynamically selects the most suitable agent at each time step. Inspired by LLM-based planning systems such as LLM-Planner (Song et al. 2023), Mic (Qiao et al. 2023), and A2Nav (Chen et al. 2023), our router leverages a large VLM model (e.g., GPT-4o (OpenAI 2024), Qwen2.5-VL (Bai et al. 2025)) in a zero-shot in-context fashion. We structure the routing process into two distinct reasoning phases:
Phase 1: Subgoal Localizer
Given the reordered subgoals $I_{\text{reorder}}=[p_{1},p_{2},\ldots,p_{m}]$, observation history $H_{t-1}$, and the sequence of previously executed subgoals $G_{t-1}=[p^{*}_{1},\dots,p^{*}_{t-1}]$, the model identifies the next subgoal $p^{*}_{t}$ to be executed for current time step $t$ and outputs the corresponding reasoning trace $r_{t}$. The output can be formalized as:
@@
Phase 2: Skill Router
At time step $t$, the skill router determines which skill-based agent $\pi^{*}_{t}\in\mathcal{S}$ is most appropriate for executing the selected subgoal $p_{t}^{*}$. , where $\mathcal{S}=\{\pi_{\text{da}},\pi_{\text{vm}},\pi_{\text{sp}},\pi_{\text{ld}},\pi_{\text{ar}}\}$ denotes the predefined set of five skill-based agents. Besides, it receives the original instruction $I$ as contextual input to capture additional linguistic cues such as verbs and spatial references. It also uses the reasoning trace $r_{t}$ from Phase 1 to enhance its understanding of the current subgoal. At each step, exactly one skill is selected, formulated as
@@
Once the appropriate skill-based agent is selected, it is invoked following the Definition 1 to predict the action for navigation at time step $t$:
@@
This routing architecture facilitates modular skill execution by leveraging natural language and perceptual inputs, with the LLM serving as a structured intermediary that bridges high-level instructions and low-level action modules.

### v2

@@ Abstract
Vision-and-Language Navigation (VLN) poses significant challenges for agents to interpret natural language instructions and navigate complex 3D environments. While recent progress has been driven by large-scale pre-training and data augmentation, current methods still struggle to generalize to unseen scenarios, particularly when complex spatial and temporal reasoning is required. In this work, we propose SkillNav, a modular framework that introduces structured, skill-based reasoning into Transformer-based VLN agents. Our method decomposes navigation into a set of interpretable atomic skills (e.g., Vertical Movement, Area and Region Identification, Stop and Pause), each handled by a specialized agent. To support targeted skill training without manual data annotation, we construct a synthetic dataset pipeline that generates diverse, linguistically natural, skill-specific instruction-trajectory pairs. We then introduce a novel training-free Vision-Language Model (VLM)-based router, which dynamically selects the most suitable agent at each time step by aligning sub-goals with visual observations and historical actions. SkillNav obtains competitive results on commonly-used benchmarks, and establishes state-of-the-art generalization to the GSA-R2R, a benchmark with novel instruction styles and unseen environments.
Codehttps://hlr.github.io/SkillNav/
@@ Introduction
Figure 1: SkillNav decomposes complex navigation instructions into atomic skills, which can be flexibly recomposed to address new environments.
Vision-and-Language Navigation (VLN) (Anderson et al. 2018; Zhang et al. 2024b) is a critical subfield of embodied AI that integrates natural language understanding, visual perception, and sequential decision-making to allow autonomous agents to navigate and interact within visual environments. With the rise of foundation models (Zhou et al. 2024a; Xiao and Zhu 2025; Li et al. 2024; Zhang et al. 2024a), VLN has seen notable progress in multimodal grounding and generalization.
@@
Despite recent advances, a key challenge in VLN lies in enabling agents to generalize reliably and interact with unseen environments and novel instructions. Previous approaches have enhanced VLN agents’ generalization ability through extensive training on large-scale synthetic instruction-trajectory pairs across varied environments (Hao et al. 2020; Chen et al. 2022a; Wang et al. 2023, 2024b). While data-driven methods improve VLN agents’ generalization, their main limitation is reliance on black-box, end-to-end models (Anderson et al. 2018; Hong et al. 2021) that tend to memorize training examples. This restricts their effectiveness in unobserved scenarios requiring deeper compositional reasoning, such as understanding diverse instructions, temporal relationships, or complex landmarks, and generalizing across a wide range of visual environments. Beyond data-driven approaches, recent work has explored zero-shot approaches leveraging Large Language Models (LLMs) for VLN tasks to improve generalization ability (Zhou, Hong, and Wu 2023; Long et al. 2024; Chen et al. 2024; Zhang et al. 2025a). Although zero-shot LLM-based agents show relatively stable performance across seen and unseen environments, they still considerably lag behind fine-tuned VLN models. Specifically, we observe a significant performance gap (approximately $36\%$ in Success Rate), primarily arising from intrinsic limitations of LLMs, including their insufficient grounding in embodied environments and imprecise alignment of linguistic instructions with specific navigational actions. This gap highlights the urgent need for methods that combine the broad generalization and compositional reasoning capabilities of LLMs with the domain-specific adaptability of fine-tuning strategies.
@@
To address these limitations, we propose SkillNav, a modular VLN framework that decomposes navigation learning into individual and reusable skills, enabling flexible re-composition and enhanced generalization in new environments (as shown in Figure 1). Unlike prior methods that treat instruction execution as an end-to-end mapping from instructions directly to actions, SkillNav explicitly captures the compositional nature of navigation tasks. Furthermore, we introduce a novel Vision-Language Model (VLM)-based router that leverages multi-modal reasoning to dynamically select the most appropriate skill at each navigation step, conditioned on the current sub-instruction, visual observation, and historical actions. SkillNav not only improves interpretability by making the decision-making processes more transparent but also facilitates robust adaptation to diverse instructions and unseen visual environments.
@@
Specifically, we build on previous research (Wang et al. 2024c), and identify a set of atomic skills required for effectively completing the VLN task. For each skill, we construct a dataset containing relevant instructions paired with corresponding visual observations, and fine-tune a dedicated agent on top of a strong VLN backbone. This process yields five specialized skill agents, each proficient in its designated capability. After obtaining these agents, we then integrate them into a unified framework to perform complex navigation tasks. Moreover, we introduce a temporal reordering module to generate chronologically ordered sub-goals, facilitating effective temporal reasoning during skill selection. Finally, we integrate a VLM-based router that dynamically identifies the next relevant sub-goal and selects the most suitable skill-based agent to execute the corresponding navigation action.
@@
SkillNav attains a strong performance on the Room-to-Room (R2R) benchmark (Anderson et al. 2018), and achieves state-of-the-art (SOTA) generalization to the GSA-R2R benchmark (Hong et al. 2025) which introduces novel instructions and diverse visual environments, including both unseen residential and non-residential settings. Additionally, we evaluate individual skill-based agents using NavNuances (Wang et al. 2024c), a dataset specifically designed for fine-grained skill evaluation. We provide comprehensive ablation studies and qualitative analysis to thoroughly assess the effectiveness of each component within our framework and justify our router design choices. Our contributions are summarized as follows:
@@
    We propose SkillNav, a modular framework that explicitly decomposes the navigation task into atomic, reusable skills, then recomposes them for execution, leveraging the specialization of fine-tuned VLN architectures together with the generalization capability of VLMs. This design significantly enhances generalization to novel instructions and visual environments.
@@
    We construct a synthetic dataset pipeline that enables skill-specific supervision without human annotation, producing diverse and linguistically natural data.
@@
    We demonstrate SOTA generalization on the challenging GSA-R2R dataset and provide a comprehensive analysis with ablation studies.
@@
Vision-and-Language Navigation Models. A wide range of methods have been proposed for addressing VLN tasks. These methods have evolved from early LSTM-based architectures (Anderson et al. 2018; Tan, Yu, and Bansal 2019) to Transformer-based models (Chen et al. 2021, 2022b; An et al. 2023) and, most recently, to Large Language Model (LLM)-based agents (Zhou, Hong, and Wu 2023; Chen et al. 2024; Lin et al. 2024; Zhou et al. 2024b; Zheng et al. 2024; Zhang et al. 2025b). A critical challenge in VLN research is enhancing the generalization capability of agents, allowing them to navigate effectively in unfamiliar environments and handle novel instructions. To enhance generalization, most existing methods utilize data-driven augmentation strategies, focusing either on augmenting visual observations (Li, Tan, and Bansal 2022; Liu et al. 2021; Li and Bansal 2023) or synthesizing additional navigation instructions (Wang et al. 2023, 2024b; Hao et al. 2020; Zhang and Kordjamshidi 2023; Zhang, Guo, and Kordjamshidi 2024). However, a fundamental limitation of purely data-driven augmentation approaches lies in their reliance on end-to-end training paradigms. Such monolithic models often memorize training examples rather than genuinely generalize, failing to fundamentally address the compositional reasoning required in novel or unseen scenarios. More recently, some approaches (Zhou, Hong, and Wu 2023; Chen et al. 2024; Long et al. 2024; Zhang et al. 2025a) have explored zero-shot navigation by heavily depending on the general reasoning capabilities of LLMs without explicit training on task-specific datasets. However, their effectiveness remains constrained by the LLMs’ inherent lack of detailed spatial understanding and precise grounding in real-world action execution. In contrast, we propose SkillNav, a modular framework that explicitly decomposes VLN tasks into reusable navigation skills. Each skill is individually fine-tuned for precise spatial grounding, while high-level reasoning and flexible skill composition leverage LLMs and VLMs, significantly improving generalization to unseen environments and varied instructions.
Skill-based MoE Systems. Mixture-of-Experts (MoE) models traditionally operate at the parameter level, distributing input across multiple expert networks to improve capacity and efficiency (Jacobs et al. 1991; Jordan and Jacobs 1994; Yuksel, Wilson, and Gader 2012). Sparsely activated MoEs (Shazeer et al. 2017; Lepikhin et al. 2021; Zhang et al. 2021; Zuo et al. 2022) further scale this idea by routing each input to a small subset of experts, making it possible to train trillion-parameter models while controlling inference cost. More recently, large language models have begun to employ skill-based MoEs at the module or LLM level, where different LLMs are specialized through fine-tuning or task profiling (Riquelme et al. 2021; Wang et al. 2024a; Dai et al. 2024; Jiang et al. 2024; Xue et al. 2024; Chen et al. 2025; Zhou et al. 2024c; Yu et al. 2025), and expert selection is performed via prompting or routing mechanisms based on task semantics. While these skill-based MoE methods focus on video understanding (Yu et al. 2025) and visual or textual question-answering (Chen et al. 2025), they largely overlook embodied tasks such as VLN. Although a recent model, SAME (Zhou et al. 2024c), introduces a state-adaptive MoE framework for VLN, this approach lacks explicit skill representations and independent spatial grounding, limiting its interpretability and extensibility. In contrast, our framework explicitly defines skill-based MoE agents for VLN tasks, employing specialized skills to significantly enhance generalization, interpretability, and extensibility.
Preliminaries
In the VLN problem setting, an agent navigates through an environment by following a natural language instruction $I$ to reach a specified target location. The environment is discretized into a connectivity graph $\mathcal{G}=(V,E)$, where $V$ denotes a non-empty set of navigable nodes, and $E$ is a set of undirected connectivity edges. At each time step $t$, the agent located at viewpoint $v_{t}$ receives a panorama represented by $n$ images, denoted as $D_{t}=\{o_{i}\}_{i=0}^{n}$. The agent is aware of a subset of views $O_{t}\subseteq D_{t}$ heading towards its navigable neighboring nodes $\mathcal{N}(v_{t})$. The local action space $A_{t}$ contains navigating to node $v\in\mathcal{N}(v_{t})$ or stopping at current node $v_{t}$.
$a^{*}_{t}=\pi(I,O_{t},M_{t}).$ (1)
where $M_{t}\subseteq\mathcal{G}$ denotes the online constructed topological map observed after $t$ steps of navigation, and $a^{*}_{t}\in A_{t}$ is the predicted action.
Methodology
@@
We propose a framework, SkillNav, for VLN that coordinates a set of atomic skill-based agents to solve navigation tasks. SkillNav enhances generalization by treating navigation as a composition of atomic skills rather than a direct language-to-action mapping. This design mirrors how humans transfer sub-skills across unfamiliar situations, preventing overfitting to specific trajectories and enabling systematic reuse of skills across environments and instruction styles. As shown in Figure 2, the framework comprises three components: a temporal reordering module for instruction decomposition, a VLM-based router for skill selection, and a set of skill-specific agents. Each agent is built upon the DUET architecture and trained with tailored synthetic data to make skill-conditioned decisions. This section introduces the proposed skill taxonomy, skill-specific synthetic dataset construction, and reasoning framework for acquiring these modular skills.
@@
We use the defined skills in NavNuances (Wang et al. 2024c) that appear to be essential for building a robust VLN agent. NavNuances provides skill categories and creates a diagnostic dataset to analyze models’ errors. However, it does not provide solutions for improving the agent skills. In this work, we extend the initially proposed skill categories and provide solutions for acquiring them by the skill-based agents. We adopt four frequently observed atomic skills from NavNuances, Direction Adjustment, Vertical Movement, Landmark Detection, and Area and Region Identification. Moreover, we find persistent challenges in temporal reasoning and stop criteria. Errors in temporal reasoning often disrupt the correct order of subgoal execution. Critical stop decisions are sometimes made too early or too late, reducing navigation success. To address these issues, we extend the skill taxonomy with two additional skills: Stop and Pause and Temporal Order Planning. In the following, we elaborate on these two new skills and their roles in navigation.
Stop and Pause captures the agent’s ability to dynamically control motion termination and temporary halting in response to visual or linguistic cues. This includes recognizing explicit stop commands (e.g., “Stop at the doorway”) or context-sensitive halts triggered by landmarks or obstacles (e.g., “Pause when you see the red sign”). The stop and pause skill emphasizes precise temporal-spatial control to ensure safe, context-aware navigation.
Temporal Order Planning reflects the agent’s capability to reason over the sequence and structure of subgoals. This includes understanding conditional immediacy (e.g., “Once you enter the hallway, turn left”), maintaining actions for a bounded duration (e.g., “Keep walking until you see the staircase”), executing forward sequential steps (e.g., “Go forward, then turn right, and finally stop”), and handling backward references to prior states (e.g., “Before turning, make sure you’re at the hallway entrance”). Effective temporal order planning involves temporal relations that guide both when and how atomic skills should be executed.
To quantify the presence and frequency of these skills in R2R (Anderson et al. 2018), we perform a keyword-based analysis of the navigation instructions as shown in Figure 4 in Appendix A. Each instruction is scanned for a curated set of indicative keywords, compiled for each skill category based on linguistic patterns observed in prior datasets and real-world navigation discourse. For instance, terms like “wait” or “stay” are used to detect Stop and Pause, while words such as “stairs” or “elevator” signal Vertical Movement. An instruction can be counted for multiple skills if it exhibits multiple relevant keywords.
@@
To enable the training of skill-specialized agents, we construct a set of synthetic datasets in which each trajectory–instruction pair is specifically designed to emphasize a single navigation skill.
We begin with a random starting node in the Matterport3D (Chang et al. 2017) environment and sample diverse navigation paths through graph traversal. For each skill, we define filtering heuristics to select trajectories where this skill is the primary factor for successful navigation. For instance, we emphasize frequent orientation changes or non-trivial turning sequences for the Direction Adjustment category. We explain detailed primary factors of skill-based trajectory generation in Appendix A. Each selected trajectory consists of a sequence of panoramic observations. Besides, we constrain trajectory length to 4–7 steps to keep the difficulty and temporal context comparable to human-annotated VLN data. The analysis of path length constraints used during trajectory generation to ensure temporal consistency and alignment with existing VLN datasets are provided in Figure 5 in Appendix B.
To generate skill-focused instruction, we feed the observation sequence of each candidate trajectory into GPT-4o (OpenAI 2024) with a structured prompt. We design the prompts such that the generated instructions preserve the general linguistic quality of real VLN datasets, including comparable sentence length, vocabulary diversity, and fluency, while emphasizing the content toward the targeted skill. This is achieved by providing GPT-4o with explicit skill-focused cues during generation, encouraging, for example, frequent references to orientation change for the Direction Adjustment skill or strong emphasis on landmark description for the Landmark Detection skill. For each skill, we synthesize $N$ such trajectory–instruction pairs, forming six separate datasets. A summary of dataset statistics is provided in Table 1.
The training of each skill-based agent is conducted in two stages. In the first stage, we fine-tune the pre-trained DUET model using the original R2R training dataset, the ScaleVLN augmentation data (Wang et al. 2023), and our Temporal Synthetic dataset to obtain a strong, skill-agnostic backbone. We provide the analysis of the effectiveness of the Temporal Order Planning agent in Appendix C. In the second stage, this backbone is further fine-tuned on a skill-specific synthetic dataset to obtain a specialized agent in the targeted skill. Following this process, we obtain five specialized skill-based agents: the Direction Adjustment agent ($\pi_{\text{da}}$), Vertical Movement agent ($\pi_{\text{vm}}$), Stop and Pause agent ($\pi_{\text{sp}}$), Landmark Detection agent ($\pi_{\text{ld}}$), and Area and Region Identification agent ($\pi_{\text{ar}}$). We denotes the predefined set of five skill-based agents as $\mathcal{S}=\{\pi_{\text{da}},\pi_{\text{vm}},\pi_{\text{sp}},\pi_{\text{ld}},\pi_{\text{ar}}\}$.
@@
After training specialized agents for different navigation skills, we build our SkillNav framework. SkillNav first employs a temporal reordering module to generate chronologically ordered execution plans. Then, we introduce a VLM-based action router to accurately identify the current subgoal and dynamically select the corresponding skill-based agent to choose the appropriate action.
Table 2: Performance comparison on R2R and GSA-R2R benchmarks. indicates large-scale data augmentation. SRDF performs best on R2R due to extensive pretraining on data that mimics R2R-style instructions; however, it struggles to generalize effectively to the GSA-R2R dataset.
@@
Mixture of Skill-based VLN
SAME (Zhou et al. 2024c) 11 $2.73$ $76$ $66$ $3.03$ $74$ $64$
SkillNav (ours) 12 $1.97$ $\mathbf{89}$ $83$ $77$ $2.53$ $83$ $78$ $70$ $\mathbf{79}$ $\mathbf{69}$ $\mathbf{72}$ $\mathbf{61}$ $\mathbf{57}$ $\mathbf{48}$
@@
The Temporal Reordering Module only takes the original natural language instruction as input. It applies the instruction reordering prompt to turn navigation instructions into a list of subgoals $I_{\text{reorder}}$. It follows the four temporal relations described in the Temporal Order Planning skill in Section Skill Taxonomy, making implicit temporal details explicit and ensuring the correct subgoal execution order. This procedure is formulated as
@@
To coordinate skill-based agents during navigation, we introduce an Action Router that dynamically selects the most suitable agent at each time step. Inspired by LLM-based planning systems such as LLM-Planner (Song et al. 2023), Mic (Qiao et al. 2023), and A2Nav (Chen et al. 2023), our router leverages a large VLM model (e.g., GPT-4o (OpenAI 2024), Qwen2.5-VL-7B-Instruct (Bai et al. 2025)) in a zero-shot in-context fashion. We structure the routing process into two distinct reasoning phases:
Phase 1: Subgoal Localizer. Given the reordered subgoals $I_{\text{reorder}}=[p_{1},p_{2},\ldots,p_{m}]$, observed history $H_{t-1}$, and the sequence of previously executed subgoals $G_{t-1}=[p^{*}_{1},\dots,p^{*}_{t-1}]$, the model identifies the next subgoal $p^{*}_{t}$ to be executed for the current time step $t$ and outputs the corresponding reasoning trace $r_{t}$, later used by the router for decision verification. The output can be formalized as:
@@
Phase 2: Skill Router. At time step $t$, the skill router determines which skill-based agent $\pi^{*}_{t}\in\mathcal{S}$ is most appropriate for executing the selected subgoal $p_{t}^{*}$. Besides, it receives the original instruction $I$ as a part of the input context to capture additional linguistic cues such as verbs and spatial references. It also uses the reasoning trace $r_{t}$ from Phase 1 to enhance its understanding of the current subgoal. At each step, exactly one skill is selected, formulated as
@@
Once the appropriate skill-based agent is selected, it is invoked by the following Equation 1 to predict the navigation action at time step $t$:
@@
Our router enables modular skill execution by integrating natural language, visual inputs, and observed history, using the Temporal Reordering LLM to bridge instructions with actionable skill modules.
@@
Table 3: Evaluation of each skill-based agent on the NavNuances benchmark across four skill categories: Direction Change (DC), Vertical Movement (VM), Landmark Recognition (LR), and Room Recognition (RR). Following the NavNuances, evaluation metrics differ across skill subsets: DC and LR are reported only with SR, VM includes SR/OSR/SPL, and $R$ provides SR/OSR. We retain this heterogeneous metric design to ensure comparability with prior work. Ident.: Identification.